# VEDA: 3D Molecular Generation via Variance-Exploding Diffusion with Annealing

### v1

@@ Abstract
Codehttps://aaai.org/example/code
Datasetshttps://aaai.org/example/datasets
Extended versionhttps://aaai.org/example/extended-version
@@ Introduction
In recent years, numerous diffusion models have been developed for 3D molecular generation. However, existing approaches have largely focused on refining their network architecture or by injecting domain-specific knowledge to guide the generation process—for instance, by developing novel equivariant transformers (liaoequiformer), shifting from direct 3D atom coordinates to internal coordinates like torsion angles (jing2022torsional), or explicitly modeling chemical bond interactions (huang2023mdm)—while overlooking the principled design of the diffusion process itself (zhang2025unraveling). A critical yet unaddressed question is how to redesign the learning objective to best match a unique characteristic of SE(3)-equivariant neural networks: their strong tendency to learn identity-like mappings, a bias stemming from their message-passing mechanism and strict geometric constraints. In diffusion models, this is often achieved through a technique called preconditioning, which provides a vital analytical framework to make the training objective more tractable (karras2022elucidating). By aligning the objective with the model’s architectural inductive biases, this principled approach can significantly improve performance and training stability. Despite its potential, this promising direction remains largely overlooked in current molecular generative models.
@@
Figure 1: An overview of the VEDA framework, detailing its training and sampling processes. During Training (top), a clean molecule $\mathcal{M}_{0}$ is perturbed via Gaussian noise for coordinates ($\mathbf{x}_{t}$) and categorical corruption for features ($\mathbf{z}_{t}$). An equivariant network $f_{\theta}$ is then trained to predict the original molecule by minimizing a combined Mean Squared Error (MSE) and Cross-Entropy (CE) loss. During Sampling (bottom), the process starts from a pure noise distribution (a Gaussian point cloud with uniform categorical features) and iteratively refines the sample over $K$ steps. Each integration step $i$ involves: (1) a noise injection from $(\mathbf{x}_{i},\mathbf{z}_{i})$ to $(\mathbf{\hat{x}}_{i},\mathbf{\hat{z}}_{i})$; (2) the preconditioned network $f_{\theta}$ uses this state to predict the final coordinates ($x_{\text{pred}}$ and category probabilities $p^{(z)}_{\theta}:=\hat{p}_{\theta}(\mathbf{z}_{0}\mid\tilde{\mathbf{z}}_{t},t)$); and (3) an update is performed by a continuous step function and a Discrete Flow Matching sampler to produce sample $\mathcal{M}_{i-1}$ in the next step. In the diagram, blue boxes represent the main network module, gray boxes are auxiliary operations, and arrows indicate the data flow.
@@
Denoising Diffusion Probabilistic Models (DDPMs) (ho2020denoising) have become a leading class of generative models which learn to reverse a forward process that incrementally corrupts data with Gaussian noise. DDPMs are extended to continuous-time settings using score-based stochastic differential equations (SDEs) (songscore), incorporating either variance-preserving or variance-exploding dynamics. Recent advances, particularly those utilizing Variance-Exploding (VE) SDEs with preconditioning techniques (karras2022elucidating), have substantially enhanced generation quality and sampling efficiency. An alternative approach, Flow Matching (lipman2023flow), directly learns the velocity field of the data distribution’s probability flow, offering a different yet highly competitive approach to generative modeling. Consistent with recent literature (gao2024diffusion), we consider flow matching a specialized formulation within the broader diffusion framework. which paper said that flow based methods can be difficult catching the distribution of confirmation generation? this does not match with the abstract.
@@
The principles of diffusion have also been extended to discrete data generation, drawing inspiration from masked language modeling (devlin2019bert). These non-autoregressive methods are a natural fit for the unordered structure of molecules, enabling de novo generation without imposing artificial sequential or ordering assumptions. D3PM (austin2021structured) established the theoretical foundations for discrete diffusion using continuous-time Markov chains. Subsequent developments, such as SEDD (lou2024discrete) and variants of Discrete Flow Matching (campbell2024generative; gat2024discrete), have further advanced the field with improved sampling strategies and scalability.
@@
We represent molecules as point clouds of chemical elements in 3D space, denoted by $\mathcal{M}=(\mathbf{x},\mathbf{z})$, where $\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{N})\in\mathbb{R}^{N\times 3}$ are the atomic coordinates of $N$ atoms, and $\mathbf{z}$ denotes non-geometric molecular attributes. $\mathbf{z}$ encodes discrete attributes such as atom types, formal charges, or bond types.
In most graph-based models, a molecule is further represented by a graph $G=(V,E)$, where nodes $V$ represent atoms and edges $E$ represent either inferred or explicitly defined interactions between atoms. These interactions can be constructed based on distance thresholds or chemical bond annotations. Some models treat bond information as part of the output (explicit bond modeling), while others infer it post hoc based on generated atomic coordinates (implicit bond inference). This distinction influences the architecture and training strategy of generative models, as detailed in later sections.
@@
The VEDA-E variant is built upon the EGNN architecture (satorras2021n_EGNN) from EDM (hoogeboom2022equivariant). In this setup, the model generates only node-centric features $\mathbf{h}=(\mathbf{h}_{1},\dots,\mathbf{h}_{N})\in\mathbb{R}^{N\times d_{h}}$, which include atom types and charges. Diffusion is performed by adding Gaussian noise to all continuous features. The chemical bond structure $\mathbf{b}$ is not explicitly modeled during generation; instead, it is inferred post hoc based on interatomic distances and chemical valence rules.
@@
For discrete features $\mathbf{z}$, we consider two variants: In VEDA-E, we apply Gaussian noise: $\mathbf{z}_{t}=\mathbf{z}_{0}+t\,\eta,\quad\eta{\sim}\mathcal{N}(0,I).$ In VEDA-S, we use time-dependent categorical corruption:
$\mathbf{z}_{t}\sim\mathrm{Cat}\Bigl[(1-m(t))\,\delta(\mathbf{z}_{t}=\mathbf{z}_{0})+\tfrac{m(t)}{S}\Bigr]$ (2)
where $S$ is the number of categories, and $m(t)=\frac{\log(t/T_{\min})}{\log(T_{\max}/T_{\min})}$ is the mask rate function that aligns the discrete corruption with the continuous noise schedule, so that at $t=T_{\max}$ all tokens are random.
Our approach uses this principled VE scheduler during training. This variance-exploding (VE) formulation is particularly suitable for modeling continuous data such as 3D molecular structures, as it allows for flexible noise schedules and accurate noise level control in log-space, injecting a massive amount of noise to reach lower energy. We believe infinite separation at $t\to\infty$ between atoms is an ideal noisy state for molecule generation over a standard Gaussian distribution as it corresponds to no chemical affiliation between atoms.
@@
Optimizing the preconditioning framework with a coordinate-space MSE loss requires the network ($F_{\theta}$) to output residuals that are uncorrelated with the input. However, our backbone (GNN or Graph Transformer) naturally outputs absolute coordinates rather than residuals, possesses a strong inductive bias toward learning identity-like mappings. As a result, $F_{\theta}$ often produces outputs highly correlated with $\mathbf{x}_{t}$, which conflicts with the residual learning objective.
To address this mismatch, we subtract an scaled identity component from the network output. Instead of forcing the network to suppress its inherent bias through the loss function, we explicitly subtract the undesired identity component from its output. Our modified denoiser is:
$D_{\theta}(\mathbf{x}_{t};t)=c_{\text{skip}}\mathbf{x}_{t}+c_{\text{out}}\left(F_{\theta}(c_{\text{in}}\mathbf{x}_{t};c_{\text{noise}})-\alpha_{t}c_{\text{in}}\mathbf{x}_{t}\right)$
The coefficient $\alpha_{t}$ is our main contribution in this section. It serves as the optimal linear predictor of the ground-truth noise $\sigma_{d}\bm{\epsilon}$ from the noisy input $\mathbf{x}_{t}$, where $\sigma_{d}$ is the standard deviation of the data distribution (karras2022elucidating). This corresponds to the Linear Minimum Mean Squared Error (LMMSE) solution, which yields the following closed-form coefficient:
$\alpha_{t}$ $=\arg\min_{\alpha}\ \mathbb{E}[\|\sigma_{d}\bm{\epsilon}-\alpha_{t}\mathbf{x}_{t}\|^{2}]$ (4)
@@
Using this optimal $\alpha_{t}$ reduces the residual correlation with the input and better aligns the training objective with the network’s inductive bias, i.e., its preference for modeling absolute coordinates.(See Appendix H for derivation.) We use the standard definitions of $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$, $c_{\text{noise}}$, and $\lambda_{\sigma}$ from karras2022elucidating.
@@
For VEDA-S, the model predicts the original discrete features $\mathbf{z}_{0}$ via a categorical distribution $\hat{p}_{\theta}(\mathbf{z}_{0}\mid\tilde{\mathbf{z}}_{t},t)$. Thisi is optimized using a KL divergence loss:
@@
We adopt the standard denoising framework where the generative process iteratively refines noisy inputs into data samples, following a predefined SDE formulation (karras2022elucidating). Conceptually, this process is analogous to simulated annealing. The noise level $t$ acts as a temperature parameter that is gradually lowered, enabling broad exploration of the energy landscape at high noise and convergence to low-energy, stable conformations as noise decreases. Our model operates under the $x$-prediction parameterization to directly estimate the clean sample from its noisy counterpart. During sampling, VEDA-E applies continuous noise injection to all features, while VEDA-S uses continuous noise injection for coordinates and discrete token refinement for categorical features. We now describe four interlocking sampling components: (1) continuous denoising with amplified noise injection, (2) discrete masked-token refinement, and (3) the proposed arcsin noise scheduler.
@@
Our sampling process employs an amplified noise injection strategy. This approach is governed by a hyperparameter, $\gamma>0$, which controls the degree of noise amplification. This process is equivalent in expectation to Gaussian smoothing on the molecular potential energy surface with a Gaussian kernel bandwidth of $\sqrt{\gamma^{2}+2\gamma}\cdot t_{i}$(See Appendix I for deviation). This pronounced smoothing suppresses local minima and surface roughness, making it easier for the sample trajectory to find the global low-energy basin and thereby improving the final energy metric (miao2015gaussian).
@@
$R^{\theta}(x_{t},j)=\omega(t)\cdot p_{\theta}(j|x_{t})+\eta_{t}\cdot p_{\theta}(j=x_{t}|x_{t})$ (10)
where $\omega(t)=\frac{\eta_{t}S(1-m(t))+\eta_{t}m(t)+m^{\prime}(t)}{m(t)}$ is the time-dependent scaling factor, and $\eta_{t}$ is based on the categorical noise hyperparameter $\eta$. To determine the optimal configuration, we performed a comprehensive grid search. The results confirmed that the variable setting from campbell2024generative, where $\eta_{t}=\frac{\eta}{m^{\prime}(t)}$, achieves the best and most robust performance, outperforming both the fixed setting ($\eta_{t}=\eta$) and the simplified DFM formulation from gat2024discrete. The transition rates are thus derived to optimally combine the model’s confidence in clean data predictions with the stability of maintaining current token assignments(see Appendices F, G for derivations and C for grid search details).
@@
We propose an arcsin-based noise scheduler that further improves the sampling process. Through empirical analysis, we observe that sampling steps corresponding to near-zero log signal-to-noise ratio (log-SNR), i.e. $\log(\text{signal variance/noise variance})$ are especially critical for final molecular structure formation. It aligns well with the distribution of $t{\sim}\exp(\mathcal{N}(\ln\sqrt{T_{\min}T_{\max}},\frac{1}{8}\ln(T_{\max}/T_{\min}))$ used during training. To leverage this observation, let $u=i/N$ is the normalized step index for a total of $N$ steps. We design an arcsin-shaped scheduler parameterized by a tunable scalar $\rho$. Our scheduler is defined as:
$w$ $=(1-\rho)\,u\;+\;\rho\,\tfrac{2}{\pi}\arcsin(\sqrt{u})$ (11)
$t^{(i)}$ $=T_{\min}\Bigl(\tfrac{T_{\max}}{T_{\min}}\Bigr)^{w}$ (12)
Where $\rho\in[0,\frac{1}{1-\frac{2}{\pi}}]$ modulates the concentration around log-SNR $\approx 0$. When $\rho=0$, the scheduler reduces to a log-uniform schedule commonly used in prior work; as $\rho$ increases, more steps cluster in the mid-range. This targeted allocation of sampling steps improves structural fidelity and overall sample quality, especially in chemically sensitive configurations. A visualization of different scheduling profiles is provided in Figure 2.
@@
EQGAT 500 0.899 0.834 1.00 1.15 8.58 6.40 11.1 0.915 0.975
JODO 500 0.963 0.879 0.77 0.83 6.01 4.74 7.04 - -
Megalodon-quick 500 0.944 0.900 0.66 0.71 5.58 3.19 5.76 - -
FlowMol2 100 0.944 0.746 1.30 1.62 15.0 17.9 24.3 0.992 1.038
SemlaFlow 100 0.969 0.920 3.10 2.06 6.05 32.3 91.0 0.273 0.358
Megalodon-flow 100 0.987 0.948 2.30 1.62 5.58 20.9 46.9 - -
VEDA-S 100 0.995 0.988 0.69 0.41 2.71 1.72 2.93 0.096 0.197
VEDA-S 50 0.968 0.966 0.86 0.60 5.10 2.99 8.38 0.236 0.355
Table 2: Performance comparison on the GEOM-DRUGS dataset. Our model, VEDA-S (bottom), is compared against denoising-based (top) and flow-based (middle) methods. Bold values indicate the best result. Metrics include Molecular Stability (MS), Validity & Connectivity (V&C), and Mean Absolute Error (MAE) for bond lengths, angles, and torsions. All metrics are lower-is-better except for MS and V&C. NFE is the Number of Function Evaluations. RMSD values for some models are missing as they were not in the original benchmark. For each model, 5000 molecules were evaluated, the full results with confidence intervals and sampling at other steps are included in Appendix B.
@@
As shown in Table 1, our model VEDA achieves state-of-the-art performance on the QM9 dataset. In the bond-implicit setting, VEDA-E achieves the highest valid and unique rate (97.9%) with only 50 NFE, significantly outperforming methods like EDM (90.9% at 1000 NFE) and GeoLDM (92.6% at 1000 NFE). In the bond-explicit setting, VEDA-S achieves nearly perfect atom stability and molecular stability, and a valid and unique score of 98.9% at just 100 NFE. Even at 30 NFE, it remains competitive or superior to most baselines, highlighting its efficiency.
Figure 4: Ablation study on key hyperparameters on GEOM-DRUGS: We report MMFF94 energy per atom and validity when varying (left) arcsin noise factor $\rho$, (middle) noise injection level $\gamma$, and (right) sampling steps. Black circles indicate selected values ($\rho$ = 2.5, $\gamma$ = 0.4, 100 steps).
@@
This superior performance is not just in final quality but also in the trade-off between sampling efficiency and generation quality, as visually demonstrated in Figure 3. The figure plots generation quality as a function of NFE. Quality is assessed by both validity and geometric stability, measured via median relaxation energy. The results are striking: VEDA-S (in red) consistently occupies the optimal region in both plots. It achieves the high efficiency of flow-based models (green) while delivering a geometric accuracy (lower $\Delta E_{relax}$) that is an order of magnitude better than all competitors.
@@
Table 3 presents an ablation study on the GEOM-DRUGS dataset. We start from the full VEDA model and progressively remove key components to assess their impact. Disabling preconditioning refers to setting $\alpha$ to be 0. It results in a small drop in validity and modest increases in RMSD and relaxation error, highlighting the stabilizing effect of preconditioning. Removing the noise injection leads to a substantial drop in all metrics, highlighting its importance in guiding the denoising process. Introducing the optimal transport (OT) alignment mechanism, as used in SemlaFlow (irwin2025semlaflow), increases validity but significantly worsens energy and RMSD. We attribute this to OT alignment violating the independence assumption between noise and original molecular coordinates, which causes the denoising direction to become overly correlated with the aligned noise. This results in a collapse of coordinate variance during sampling and degrades overall sample quality. These results demonstrate that each module in VEDA provides complementary improvements to both accuracy and validity.
@@
Table 3: Ablation study on GEOM-DRUGS. Each variant enables (✓) or disables (✗) key components of VEDA: OT alignment (OT), noise injection (Noise), and preconditioning (Pre). V&C ↑ and M.S.↑ are reported as percentage-based metrics.

### v2

@@ Abstract
Codehttps://github.com/peiningzhang/VEDA
Extended versionhttps://arxiv.org/abs/2511.09568
@@ Introduction
In recent years, numerous diffusion models have been developed for 3D molecular generation or 3D molecular conformation generation. However, existing approaches have largely focused on refining their network architecture or by injecting domain-specific knowledge to guide the generation process—for instance, by developing novel equivariant transformers (liaoequiformer), shifting from direct 3D atom coordinates to internal coordinates like torsion angles (jing2022torsional), or explicitly modeling chemical bond interactions (huang2023mdm)—while overlooking the principled design of the diffusion process itself (zhang2025unraveling). A critical yet unaddressed question is how to redesign the learning objective to best match a unique characteristic of SE(3)-equivariant neural networks: their strong tendency to learn identity-like mappings, a bias stemming from their message-passing mechanism and strict geometric constraints. In diffusion models, this is often achieved through a technique called preconditioning, which provides a vital analytical framework to make the training objective more tractable (karras2022elucidating). By aligning the objective with the model’s architectural inductive biases, this principled approach can significantly improve performance and training stability. Despite its potential, this promising direction remains largely overlooked in current molecular generative models.
@@
Figure 1: An overview of the VEDA framework, detailing its training and sampling processes. During Training (top), a clean molecule $\mathcal{M}_{0}$ is perturbed via Gaussian noise for coordinates ($\mathbf{x}_{t}$) and categorical corruption for features ($\mathbf{z}_{t}$), defined by Eq. 1 and Eq. 3. The equivariant network $f_{\theta}$ is then trained to predict the original molecule by minimizing the combined Mean Squared Error (MSE) and Cross-Entropy (CE) loss in Eq. 12. During Sampling (bottom), the process starts from a pure noise distribution (a Gaussian point cloud with uniform categorical features) and iteratively refines the sample over $K$ steps. Each integration step $i$ involves: (1) a noise injection from $(\mathbf{x}_{i},\mathbf{z}_{i})$ to $(\mathbf{\hat{x}}_{i},\mathbf{\hat{z}}_{i})$; (2) the network $f_{\theta}$ is applied to both coordinates and features, with preconditioning affecting only the coordinate predictions; it outputs $x_{\text{pred}}$ and the category probabilities $p^{(z)}_{\theta}$ (Eq. 6); and (3) an update combining a continuous Euler step and a Discrete Flow Matching sampler to obtain $\mathcal{M}_{i-1}$, formally given in Eq. 13 and Eq. 14. In the diagram, blue boxes represent the main network module, gray boxes are auxiliary operations, and arrows indicate the data flow.
@@
Denoising Diffusion Probabilistic Models (DDPMs) (ho2020denoising) have become a leading class of generative models which learn to reverse a forward process that incrementally corrupts data with Gaussian noise. DDPMs are extended to continuous-time settings using score-based stochastic differential equations (SDEs) (songscore), incorporating either variance-preserving or variance-exploding dynamics. Recent advances, particularly those utilizing Variance-Exploding (VE) SDEs with preconditioning techniques (karras2022elucidating), have substantially enhanced generation quality and sampling efficiency. An alternative approach, Flow Matching (lipman2023flow), directly learns the velocity field of the data distribution’s probability flow, offering a different yet highly competitive approach to generative modeling. Consistent with recent literature (gao2024diffusion), we consider flow matching a specialized formulation within the broader diffusion framework.
@@
The principles of diffusion have also been extended to discrete data generation, drawing inspiration from masked language modeling (devlin2019bert). These non-autoregressive methods are a natural fit for the unordered structure of molecules, enabling de novo generation without imposing artificial sequential or ordering assumptions. D3PM (austin2021structured) established the theoretical foundations for discrete diffusion using continuous-time Markov chains. Subsequent developments of variants of Discrete Flow Matching (campbell2024generative; gat2024discrete) have further advanced the field with improved sampling strategies and scalability.
@@
We represent molecules as point clouds of chemical elements in 3D space, denoted by $\mathcal{M}=(\mathbf{x},\mathbf{z})$, where $\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{N})\in\mathbb{R}^{N\times 3}$ are the atomic coordinates of $N$ atoms, and $\mathbf{z}$ denotes non-geometric molecular attributes. To keep notation unified, we decompose $\mathbf{z}=(\mathbf{h},\mathbf{b})$, where $\mathbf{h}$ includes atom-level properties (e.g., atom types, charges) and $\mathbf{b}$ denotes pairwise bond information.
In most graph-based models, a molecule is further represented by a graph $G=(V,E)$, where nodes $V$ represent atoms and edges $E$ represent either inferred or explicitly defined interactions between atoms. These interactions can be constructed based on distance thresholds or chemical bond annotations. Some models treat bond information as part of the output (explicit bond modeling), while others infer it post hoc based on generated atomic coordinates (implicit bond inference). This distinction influences the architecture and training strategy of generative models, as detailed later.
@@
The VEDA-E variant is built upon the EGNN architecture (satorras2021n_EGNN) from EDM (hoogeboom2022equivariant). In this setup, the model generates only node-centric features $\mathbf{h}=(\mathbf{h}_{1},\dots,\mathbf{h}_{N})\in\mathbb{R}^{N\times d_{h}}$, which include atom types and charges. Diffusion is performed by adding Gaussian noise to all continuous features. The chemical bond structure $\mathbf{b}$ is not explicitly modeled during generation; instead, it is inferred post hoc based on interatomic distances and chemical valence rules, a standard practice in EDM-based molecule generation.
@@
For discrete features $\mathbf{z}$, we consider two variants.
In VEDA-E, Gaussian noise is applied:
$\mathbf{z}_{t}=\mathbf{z}_{0}+t\,\eta,\quad\eta\sim\mathcal{N}(0,I).$ (2)
In VEDA-S, we use time-dependent categorical corruption:
$\mathbf{z}_{t}\sim\mathrm{Cat}\!\left[(1-m(t))\,\delta(\mathbf{z}_{t}=\mathbf{z}_{0})+\frac{m(t)}{S}\right],$ (3)
where $S$ is the number of categories, and the masking rate function $m(t)=(\ln t-\ln T_{\min})/(\ln T_{\max}-\ln T_{\min})$ aligns the discrete corruption schedule with the continuous noise schedule, reaching uniform distribution at $t=T_{\max}$.
Our approach uses this principled VE scheduler during training. This variance-exploding (VE) formulation is particularly suitable for modeling continuous data such as 3D molecular structures, as it enables flexible noise schedules and accurate noise level control in log-space, injecting a massive amount of noise to reach lower energy. We believe infinite separation at $t\to\infty$ between atoms is an ideal noisy state for molecule generation, superior to a standard Gaussian prior, as it corresponds to no chemical interactions between atoms.
@@
$F_{\text{target}}$ $=\frac{1}{c_{\text{out}}}(\mathbf{x}_{0}-c_{\text{skip}}(\mathbf{x}_{0}+t\,\epsilon))$ (5)
Optimizing the preconditioning framework with a coordinate-space MSE loss requires the network ($F_{\theta}$) to output residuals that are uncorrelated with the input. However, our backbone (GNN or Graph Transformer) outputs coordinates via a residual connection design (he2016deep), which causes $F_{\theta}$ to produce outputs highly correlated with $\mathbf{x}_{t}$, conflicting with the residual learning objective.
To address this mismatch, we subtract a scaled identity component from the network output. Instead of forcing the network to suppress its inherent bias through the loss function, we explicitly subtract the undesired identity component from its output. Our modified denoiser is:
$D_{\theta}(\mathbf{x}_{t};t)=c_{\text{skip}}\mathbf{x}_{t}+c_{\text{out}}\left(F_{\theta}(c_{\text{in}}\mathbf{x}_{t};c_{\text{noise}})-\alpha_{t}\,c_{\text{in}}\,\mathbf{x}_{t}\right)$ (6)
The coefficient $\alpha_{t}$ is our main contribution in this section. It serves as the optimal linear predictor of the ground-truth noise $\frac{\sigma_{d}\bm{\epsilon}}{\sqrt{\sigma_{d}^{2}+t^{2}}}$ from $F_{\text{target}}$, where $\sigma_{d}$ is the standard deviation of the data distribution (karras2022elucidating). This corresponds to the Linear Minimum Mean Squared Error (LMMSE) solution, which yields the following closed-form coefficient:
$\alpha_{t}$ $=\arg\min_{\alpha}\ \mathbb{E}[\|\frac{\sigma_{d}\bm{\epsilon}}{\sqrt{\sigma_{d}^{2}+t^{2}}}-\alpha_{t}\,c_{\text{in}}\,\mathbf{x}_{t}\|^{2}]$ (7)
$=\arg\min_{\alpha}\ \mathbb{E}[\|\sigma_{d}\bm{\epsilon}-\alpha_{t}\mathbf{x}_{t}\|^{2}]$ (8)
@@
Using this optimal $\alpha_{t}$ reduces the residual correlation with the input and better aligns the training objective with the network’s inductive bias, i.e., its preference for modeling absolute coordinates.(See Appendix I for derivation.) We use the standard definitions of $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$, $c_{\text{noise}}$, and $\lambda_{\sigma}$ from karras2022elucidating.
@@
For VEDA-S, the model predicts the original discrete features $\mathbf{z}_{0}$ via a categorical distribution $\hat{p}_{\theta}(\mathbf{z}_{0}\mid\tilde{\mathbf{z}}_{t},t)$. This is optimized using a KL divergence loss:
@@
We adopt the standard denoising framework where the generative process iteratively refines noisy inputs into data samples, following a predefined SDE formulation (karras2022elucidating). Conceptually, this process is analogous to simulated annealing. The noise level $t$ acts as a temperature parameter that is gradually lowered, enabling broad exploration of the energy landscape at high noise and convergence to low-energy, stable conformations as noise decreases. Our model operates under the $x$-prediction parameterization to directly estimate the clean sample from its noisy counterpart. During sampling, VEDA-E applies continuous noise injection to all features, while VEDA-S uses continuous noise injection for coordinates and discrete token refinement for categorical features. We now describe three interlocking sampling components: (1) continuous denoising with amplified noise injection, (2) discrete masked-token refinement, and (3) the proposed arcsin noise scheduler.
@@
Our sampling process employs an amplified noise injection strategy. This approach is governed by a hyperparameter, $\gamma>0$, which controls the degree of noise amplification. This process is equivalent in expectation to Gaussian smoothing on the molecular potential energy surface with a Gaussian kernel bandwidth of $\sqrt{\gamma^{2}+2\gamma}\cdot t_{i}$ (See Appendix J for deviation) . This pronounced smoothing suppresses local minima and surface roughness, making it easier for the sample trajectory to find the global low-energy basin and thereby improving the final energy metric (miao2015gaussian).
@@
$R_{\theta}(z_{t},j)=\omega(t)\cdot p_{\theta}(z_{0}=j|z_{t})+\eta_{t}\cdot p_{\theta}(z_{0}=z_{t}|z_{t})$ (14)
where $\omega(t)=\frac{\eta_{t}S(1-m(t))+\eta_{t}m(t)+m^{\prime}(t)}{m(t)}$ is the time-dependent scaling factor, and $\eta_{t}$ is based on the categorical noise hyperparameter $\eta$. To determine the optimal configuration, we performed a comprehensive grid search. The results confirmed that the variable setting from campbell2024generative, where $\eta_{t}=\frac{\eta}{m^{\prime}(t)}$, achieves the best and most robust performance, outperforming both the fixed setting ($\eta_{t}=\eta$) and the simplified DFM formulation from gat2024discrete. The transition rates are thus derived to optimally combine the model’s confidence in clean data predictions with the stability of maintaining current token assignments(see Appendices G, H for derivations and C for grid search details).
@@
We propose an arcsin-based noise scheduler that further improves the sampling process. Through empirical analysis, we observe that sampling steps corresponding to near-zero log signal-to-noise ratio (log-SNR), i.e. $\log(\text{signal variance/noise variance})$ are especially critical for final molecular structure formation. It aligns well with the distribution of $t\sim\operatorname{LogNormal}\left(\ln\sqrt{T_{\min}T_{\max}},[\frac{1}{8}\ln(T_{\max}/T_{\min})]^{2}\right)$ used during training. To leverage this observation, let $u=i/N$ is the normalized step index for a total of $N$ steps. We design an arcsin-shaped scheduler parameterized by a tunable scalar $\rho$. Our scheduler is defined as:
$w(u)$ $=(1-\rho)\,u\;+\;\rho\,\tfrac{2}{\pi}\arcsin(\sqrt{u})$ (15)
$t^{(i)}$ $=T_{\min}\Bigl(\tfrac{T_{\max}}{T_{\min}}\Bigr)^{w(u)}$ (16)
Where $\rho\in[0,\tfrac{\pi}{\pi-2}]$ modulates the concentration around log-SNR $\approx 0$. When $\rho=0$, the scheduler reduces to a log-uniform schedule commonly used in prior work; as $\rho$ increases, more steps cluster in the mid-range. This targeted allocation of sampling steps improves structural fidelity and overall sample quality, especially in chemically sensitive configurations. As shown in Figure 2, the arcsin scheduler ($\rho{=}2$) closely matches the log-normal training distribution, further supporting its design rationale.
@@
EQGAT 500 3468 0.899 0.834 1.00 1.15 8.58 6.40 11.1 0.915 0.975
JODO 500 673 0.963 0.879 0.77 0.83 6.01 4.74 7.04 - -
Megalodon-quick 500 - 0.944 0.900 0.66 0.71 5.58 3.19 5.76 - -
FlowMol2 100 126.3 0.944 0.746 1.30 1.62 15.0 17.9 24.3 0.992 1.038
SemlaFlow 100 150.3 0.969 0.920 3.10 2.06 6.05 32.3 91.0 0.273 0.358
Megalodon-flow 100 - 0.987 0.948 2.30 1.62 5.58 20.9 46.9 - -
VEDA-S 100 164.8 0.995 0.988 0.69 0.41 2.71 1.72 2.93 0.096 0.197
VEDA-S 50 84.1 0.968 0.966 0.86 0.60 5.10 2.99 8.38 0.236 0.355
Table 2: Performance comparison on the GEOM-DRUGS dataset. Our model, VEDA-S (bottom), is compared against denoising-based (top) and flow-based (middle) methods. Bold values indicate the best result. Metrics include Molecular Stability (MS), Validity & Connectivity (V&C), and Mean Absolute Error (MAE) for bond lengths, angles, and torsions. All metrics are lower-is-better except for MS and V&C. Time (s) reports the average wall-clock time for generating one molecule. NFE is the Number of Function Evaluations. RMSD values for some models are missing as they were not in the original benchmark. For each model, 5000 molecules were evaluated, the full results with confidence intervals and sampling at other steps are included in Appendix B.
@@
As shown in Table 1, our model VEDA achieves state-of-the-art performance on the QM9 dataset. In the bond-implicit setting, VEDA-E achieves the highest valid and unique rate (97.9%) with only 50 NFE, significantly outperforming methods like EDM (90.9% at 1000 NFE) and GeoLDM (92.6% at 1000 NFE). In the bond-explicit setting, VEDA-S achieves nearly perfect atom stability and molecular stability, and a valid and unique score of 98.9% at just 100 NFE. Even at 30 NFE, it remains competitive with most baselines, highlighting its efficiency.
Figure 4: Ablation study of key hyperparameters on GEOM-DRUGS: We report MMFF94 energy and validity when varying (left) arcsin noise factor $\rho$, (middle) noise injection level $\gamma$, and (right) sampling steps. Black circles indicate selected values ($\rho{=}2.5$, $\gamma{=}0.4$, 100 steps).
@@
This superior performance is not just in final quality but also in the trade-off between sampling efficiency and generation quality, as shown in Figure 3. The figure plots generation quality as a function of NFE. Quality is assessed by both validity and geometric stability, measured via median relaxation energy. The results are striking: VEDA-S (in red) consistently occupies the optimal region in both plots. It achieves the high efficiency of flow-based models (green) while delivering a geometric accuracy (lower $\Delta E_{relax}$) that is an order of magnitude better than all competitors.
@@
Table 3 presents an ablation study on the GEOM-DRUGS dataset. We start from the full VEDA model and progressively remove key components to assess their impact. Disabling preconditioning refers to setting $\alpha_{t}$ to be 0. It results in a small drop in validity and modest increases in RMSD and relaxation error, showing the stabilizing effect of preconditioning. Removing the noise injection leads to a substantial drop in all metrics, highlighting its importance in guiding the denoising process. Introducing the optimal transport (OT) alignment mechanism, as used in SemlaFlow (irwin2025semlaflow), increases validity but significantly worsens energy and RMSD. We attribute this to OT alignment violating the independence assumption between noise and original molecular coordinates, which causes the denoising direction to become overly correlated with the aligned noise. This results in a collapse of coordinate variance during sampling and degrades overall sample quality. These results show that each module in VEDA provides complementary improvements to accuracy and validity.
@@
Table 3: Ablation study on GEOM-DRUGS. Each variant enables (✓) or disables (✗) key components of VEDA: OT alignment (OT), noise injection (Noise), and preconditioning (Pre). V&C ↑ and M.S.↑ are reported as percentage-based metrics. The last two rows share the same component configuration and differ only in their time parameterization (VE vs. FM).